# Approximate solutions to the shrinking core model

### v1

@@ 2.2 The standard Pseudo-Steady State model
The simplest manner to deal with the SCM system is to apply the Pseudo-Steady State (PSS) approximation. This is based on setting $\mathrm{Da}\ll 1$, so that the time derivative may be neglected in the diffusion equation. In terms of chemical reactions it is explained in [25] that the PSS hypothesis typically holds when a process involves two or more distinct time scales. In the present case diffusion is much more rapid than the shrinkage rate of the core. Henceforth we will take $c_{R}(t)=1$, assuming there is a large supply of reactant outside of the sphere.
@@
$c=1+B-\frac{B}{r}\,,\quad\mbox{where}\quad B=\frac{T_{m}s^{2}}{1+T_{m}s-T_{m}s^{2}}=\frac{s^{2}}{T_{m}^{-1}+s-s^{2}}>0\,.$ (32)
@@
$\frac{\text{d}s}{\text{d}t}=\frac{-1}{T_{m}^{-1}+s-s^{2}}\,<0\,.$ (33)
@@
$t_{f}=\frac{1}{6}+T_{m}^{-1}\,.$ (35)
Various versions of the PSS solution may be found throughout the literature. Based on physical arguments Levenspiel [17] defines three forms of model, controlled by: diffusion through the gas film; diffusion through the ash layer; chemical reaction. The first of these we neglect, assuming the fluid concentration surrounding the sphere remains constant throughout the process. The second refers to diffusion through the $\beta$-phase controlling the process. The third is controlled by the reaction at the interface. These may be thought of in terms of the parameter $T_{m}=\tau_{D}/\tau_{k}$. If kinetics controls the process, i.e. the reaction is much slower than the diffusion rate, $\tau_{k}\gg\tau_{D}$, then the Kinetic Controlled result of Levenspiel [17, Ch. 25.2] comes from setting $T_{m}\ll 1$ in (34) so that
$s\approx 1-T_{m}t\,.$ (36)
The Diffusion Controlled result, $\tau_{D}\gg\tau_{k}$, follows from the large $T_{m}$ limit
@@
These two limiting cases, widely accepted by the community, must be treated with caution. Firstly, the PSS approximation is based on the assumption $\mathrm{Da}={M_{g}(c_{R,0}^{*}-c_{eq}^{*})}/{\rho_{\beta}}\ll 1$. This is often quoted in various guises. Thinking of $\mathrm{Da}$ as the ratio of fluid density to solid, it is clear that for a gas-solid system $\mathrm{Da}\ll 1$ and the PSS is almost guaranteed to hold for a wide range of times and values of $T_{m}$. For a liquid-solid system the densities may be similar and there is no guarantee $\mathrm{Da}\ll 1$. The applicability of the PSS model to liquid-solid systems has been previously discussed in [8, 23, 19].
If we neglect only terms of order $\mathrm{Da}$ we must be aware that retained terms are of a greater magnitude. Specifically, the reduction requires $T_{m}\gg\mathrm{Da}$ if terms of order $T_{m}={R^{*}k}/{D}$ are retained. The kinetic controlled result, Eq. (36), therefore only holds for $\mathrm{Da}\ll T_{m}\ll 1$ and sufficiently large times, such that $T_{m}t={\cal O}\left(1\right)$ or, in dimensional form, $t^{*}={\cal O}\left(\tau_{g}/T_{m}\right)$ where $\tau_{g}/T_{m}=\rho_{\beta}R^{*}/(kM_{g}(c_{R,0}^{*}-c_{eq}^{*}))$. Since $s={\cal O}\left(1\right)$, the Diffusion Limited case holds for all time.
@@ 3 Approximate analytical solutions
In this section, we derive approximate analytical solutions to the system (26)–(27). First, we obtain a correction to the pseudo-steady-state (PSS) solution by constructing a perturbation expansion of the concentration in powers of $\mathrm{Da}$. We then analyze the behaviour of the system at early times and derive explicit expressions for both the concentration and the position of the reacting front in the small-time limit.
@@ 3.1 Perturbation solution for small $\mathrm{Da}$
$+\left.\frac{12\mathrm{T}_{m}^{-2}}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\left(\text{Arctanh}\left(\frac{1-2s}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\right)-\text{Arctanh}\left(\frac{1}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\right)\right)\right]\,.$
@@
$t_{f}=\frac{1}{6}\left[1+\mathrm{Da}(1-6\mathrm{T}_{m}^{-1})\right]+T_{m}^{-1}\,,$ (49)
@@
Similar to the PSS case, an expression for the kinetically controlled regime can be obtained by assuming $\mathrm{T}_{m}\ll 1$ in the perturbation solution (3.1). This gives
$s\approx 1-\frac{\mathrm{T}_{m}}{1-\mathrm{Da}}t\,.$ (50)
@@ 3.2 Small time solution
As discussed in Section 2.2, the PSS model is based on neglecting the time derivative in the diffusion equation, based on the assumption $\mathrm{Da}\ll 1$. If we consider a very early time-scale $\tau={\cal O}\left(\mathrm{Da}\right)$ or smaller then the PSS model no longer holds. In this section, we develop an approximate analytical solution valid in the small-time regime, enabling us to better understand the system’s initial transient behavior and provide an initial condition for a numerical solution.
@@
As only a short period of time has elapsed, the interface is not far away from $R$, the radius of the particle. Therefore, it is useful to propose an ansatz as
@@
This solution coincides with the kinetic controlled result (36).
@@
An interesting point is that although the PSS does not hold at small times, the small time behaviour matches the kinetic controlled solution (even in diffusion dominated situations). This means that, even though the kinetic solution does not hold for early times, according to the restrictions of the PSS approximation, it still captures the correct behaviour. This move away from the diffusion form may be inferred from the solution (34) by setting $s=1-\varepsilon f(t)$, where $\varepsilon\ll 1$
$t$ $=\frac{s^{3}-1}{3}-\frac{s^{2}-1}{2}-T_{m}^{-1}(s-1)=-\frac{3\varepsilon f}{3}-\frac{2\varepsilon f}{2}-T_{m}^{-1}(\varepsilon f)+{\cal O}\left(\varepsilon^{2}\right)\,$ (68)
@@ 4 Numerical solution
The semi-implicit scheme (4) can be formulated as a matrix linear system which can be solved by inverting the matrix of the system at each time-step. All numerical solutions were obtained in a uniform mesh of $10^{3}$ points, using a time-step size of $\Delta t=4\times 10^{-6}$. The implementation was carried out using a custom Matlab code.
@@ 5 Results and discussion
In this section, we first compare the approximate analytical solutions presented in Section 3 with the numerical results from Section 4, focusing on the solution behavior across different limits of the dimensionless parameters $\mathrm{Da}$ and $\mathrm{T}_{m}$. We then demonstrate how the analytical solutions of the SCM found can be used to fit experimental data for the purpose of estimating unknown physical parameters of the process.
@@ 5.1 Discussion of solutions and limiting behaviours
In typical SCM-based experiments, measurements are used to determine the reacted fraction $X(t)$, which, as shown in (31), depends on the position of the reaction front $s(t)$. Thus, the evolution of $X(t)$ is directly governed by the behavior of $s(t)$.
In Figure 2 we show the evolution of $s(t)$ for various combinations of $\mathrm{Da}=0.01$, 0.10.10.1, 111, and $\mathrm{T}_{m}=0.1$, 111, 101010, as predicted by the numerical solution, the pseudo-steady-state (PSS) approximation (34), and the first-order perturbation solution (3.1). The values of $\mathrm{Da}$ are kept below 1, as this range reflects the most physically realistic scenarios, corresponding to typical fluid-to-solid density ratios, as discussed in Section 2.1. Then, the values $\mathrm{T}_{m}=0.1$ to $\mathrm{T}_{m}=10$ are chosen to study the diffusion and reaction controlled regimes, respectively, which are the ones typically discussed in the literature.
@@
Focusing on the comparison between the numerical and approximate analytical solutions, we observe that the discrepancies increase with larger values of $\mathrm{Da}$. This behavior is expected: the PSS solution is derived in the limit $\mathrm{Da}=0$, and the first-order perturbation solution is based on an expansion for $\mathrm{Da}\ll 1$. Therefore, the smaller the value of $\mathrm{Da}$, the closer the approximate solutions are to the full numerical solution, which retains all terms in the model. This trend is especially noticeable in the cases with $\mathrm{T}_{m}=10$ (right panels of Figure 2).
@@
An interesting feature is the distinct change in the behavior of $s(t)$ around $\mathrm{T}_{m}=1$, which marks the transition between the diffusion-controlled ($\mathrm{T}_{m}\gg 1$) and reaction-controlled ($\mathrm{T}_{m}\ll 1$) regimes. In particular, the reaction-controlled regime exhibits a virtually linear evolution of $s(t)$ (left panels of Figure 2), consistent with the linear expressions obtained in (36) and (50) for the PSS and perturbation solutions, respectively, in the limit $\mathrm{T}_{m}\ll 1$. The trends observed for large $\mathrm{T}_{m}$ are characteristic of the classical Stefan problem in spherical geometry (see, e.g., [11, 29, 30]). This is expected, as setting $\mathrm{T}_{m}=0$ in equations (26)–(29), previously rewriting (29) in terms of the gradient of $c$ using (28), reduces the SCM to the classical one-phase Stefan problem for a sphere.
@@
The case with $\mathrm{Da}=0.1$, $\mathrm{T}_{m}=0.1$, corresponding to the reaction-limited regime, is characterized by concentrations remaining close to 1 throughout the domain. This reflects the fact that diffusion is much faster than reaction, allowing the concentration near the moving boundary to rapidly equilibrate with the external concentration due to efficient mass transport. The inset plot for $\mathrm{Da}=0.1$, $\mathrm{T}_{m}=0.1$ provides a close-up view of the concentration profiles near $c\approx 1$, revealing shapes qualitatively similar to those in the diffusion-limited regime, though with concentration variations that are an order of magnitude smaller. Finally, we note the excellent agreement between the numerical solution and the first-order perturbation approximation, with only minor discrepancies appearing near the end of the process in the diffusion-limited case.
@@ 5.2 Comparison to synthetic data
Experimental studies often employ fitting procedures to determine unknown physical parameters of the process. In solid-gas reaction experiments using the SCM, typical unknowns include the reaction rate constant $k$, the diffusion coefficient of the ash layer $D$ or the equilibrium concentration $c_{eq}^{*}$. The experimental data commonly consist of the reacted fraction $X$ measured as a function of time, $t^{*}$. Therefore, the fitting process generally involves estimating the unknown parameters by fitting datasets of the form $(t_{\text{exp},i},X_{\text{exp},i})_{i\in\{1,\dots,N\}}$. In this section, we show how the small time limit (61) and the perturbation solution (3.1) can be used to determine the parameters $k$, $D$ and $c_{eq}^{*}$ from an experimental dataset. In addition, the PSS approximation (34) will also be tested as it is widely used in the literature.
@@
The fitting process is done in two steps. The first, involves finding a reference value of the parameter group $\tau_{\text{eff}}\equiv\tau/\mathrm{T}_{m}$ using the dimensional form of the small time solution (61), i.e. $s(t^{*})=1-(\mathrm{T}_{m}/\tau)\,t^{*}=1-(t^{*}/\tau_{\text{eff}})$. Since the small time limits are short-lived, we take the first two points of the dataset to compute the slope $1/\tau_{\text{eff}}$ of the straight line connecting them. The second, involves using (3.1) to find the values of $T_{m}$ and $\mathrm{Da}$, that best minimize the quadratic error
@@
In Figure 4, we present the predictions of $s(t^{*})$ and $X(t^{*})$ obtained from the fitting procedure using both the first-order perturbation and PSS solutions, plotted alongside the synthetic data. For the dataset corresponding to $\mathrm{Da}=0.1$, both the first-order perturbation and PSS solutions closely match the data, and the improvement offered by the perturbation solution is not visually evident. In contrast, for the second dataset ($\mathrm{Da}=1$), the first-order perturbation solution provides a visibly better fit than the PSS. This result is expected, since the first-order correction is of order $\mathcal{O}(\mathrm{Da})$; as $\mathrm{Da}$ increases, the correction captures the behavior of $s(t^{*})$ and $X(t^{*})$ more accurately than the PSS solution, which assumes $\mathrm{Da}=0$.
The last step requires translating the results of the fitting procedure to values of physical process parameters. As mentioned previously, in solid-gas reaction systems, the most common unknowns are $k$, $D$ and $c_{eq}^{*}$. Recalling the definitions of the fitting parameters in terms of the original physical quantities
$\mathrm{Da}=\frac{M_{g}(c_{R}^{*}-c_{eq}^{*})}{\rho_{\beta}},\quad\mathrm{T}_{m}=\frac{R^{*}k}{D},\quad\tau=\mathrm{T}_{m}\,\tau_{\text{eff}}=\frac{\rho_{\beta}R^{*2}}{M_{g}(c_{R}^{*}-c_{eq}^{*})D}\,,$
@@
$c_{eq}^{*}=c_{R}^{*}+\frac{\mathrm{Da}\rho_{\beta}}{M_{g}}\,,\quad D=\frac{\rho_{\beta}R^{*2}}{M_{g}(c_{R}^{*}-c_{eq}^{*})\mathrm{T}_{m}\,\tau_{\text{eff}}}\,,\quad k=\frac{\mathrm{T}_{m}D}{R^{*}}\,,$
if the 1st order perturbation solution is used to fit the data. Although, as demonstrated, the PSS model tends to produce larger errors compared to the first-order correction during fitting, it may still yield approximate values for $k$ and $D$, provided that $c_{eq}^{*}$ is either known or negligible relative to $c_{R}^{*}$. In such cases, $c_{eq}^{*}$ can be omitted from the expressions for $\tau$ and $\mathrm{Da}$.

### v2

@@ 2.2 The standard Pseudo-Steady State model
The simplest manner to deal with the SCM system is to apply the pseudo-steady-state approximation. This is based on setting $\mathrm{Da}\ll 1$, so that the time derivative may be neglected in the diffusion equation. In terms of chemical reactions it is explained in [36] that the PSS hypothesis typically holds when a process involves two or more distinct time scales. In the present case diffusion is much more rapid than the shrinkage rate of the core. Henceforth we will take $c_{R}(t)=1$, assuming there is a large supply of reactant outside of the sphere or it is constantly introduced into the system.
@@
$c=1+B-\frac{B}{r}\,,\quad\mbox{where}\quad B=\frac{\mathrm{T}_{m}s^{2}}{1+\mathrm{T}_{m}s-\mathrm{T}_{m}s^{2}}=\frac{s^{2}}{\mathrm{T}_{m}^{-1}+s-s^{2}}>0\,.$ (32)
@@
$\frac{\text{d}s}{\text{d}t}=\frac{-1}{\mathrm{T}_{m}^{-1}+s-s^{2}}\,<0\,.$ (33)
@@
$t_{f}=\frac{1}{6}+\mathrm{T}_{m}^{-1}\,.$ (35)
Various versions of the PSS solution may be found throughout the literature. Based on physical arguments Levenspiel [28] defines three forms of model, controlled by: diffusion through the gas film; diffusion through the ash layer; chemical reaction. The first of these we neglect, assuming the fluid concentration surrounding the sphere remains constant throughout the process. The second refers to diffusion through the $\beta$-phase controlling the process. The third is controlled by the reaction at the interface. These may be thought of in terms of the parameter $\mathrm{T}_{m}=\tau_{D}/\tau_{k}$. If kinetics controls the process, i.e. the reaction is much slower than the diffusion rate, $\tau_{k}\gg\tau_{D}$, then the Kinetic-Controlled result of Levenspiel [28, Ch. 25.2] comes from setting $\mathrm{T}_{m}\ll 1$ in (34) so that
$s\approx 1-\mathrm{T}_{m}t\,.$ (36)
The Diffusion-Controlled result, $\tau_{D}\gg\tau_{k}$, follows from the large $\mathrm{T}_{m}$ limit
@@
These two limiting cases, widely accepted by the community, must be treated with caution. Firstly, the PSS approximation is based on the assumption $\mathrm{Da}={M_{g}(c_{R,0}^{*}-c_{\rm eq}^{*})}/{\rho_{\beta}}\ll 1$. This is often quoted in various guises. Thinking of $\mathrm{Da}$ as the ratio of fluid density to solid, it is clear that for a gas-solid system $\mathrm{Da}\ll 1$ and the PSS is almost guaranteed to hold for a wide range of times and values of $\mathrm{T}_{m}$. For a liquid-solid system the densities may be similar and there is no guarantee $\mathrm{Da}\ll 1$. The applicability of the PSS model to liquid-solid systems has been previously discussed in [19, 34, 30].
If we neglect only terms of order $\mathrm{Da}$ we must be aware that retained terms are of a greater magnitude. Specifically, the reduction requires $\mathrm{T}_{m}\gg\mathrm{Da}$ if terms of order $\mathrm{T}_{m}={R^{*}k}/{D}$ are retained. The kinetic-controlled result, Eq. (36), therefore only holds for $\mathrm{Da}\ll\mathrm{T}_{m}\ll 1$ and sufficiently large times, such that $\mathrm{T}_{m}t=\mathcal{O}(1)$ or, in dimensional form, $t^{*}=\mathcal{O}(\tau_{g}/\mathrm{T}_{m})$ where $\tau_{g}/\mathrm{T}_{m}=\rho_{\beta}R^{*}/(kM_{g}(c_{R,0}^{*}-c_{\rm eq}^{*}))$. Since $s=\mathcal{O}(1)$, the diffusion limited case holds for all time.
@@ 3 Approximate analytical solutions
In this section, we derive approximate analytical solutions to the system (26)–(27). First, we obtain a correction to the pseudo-steady-state solution by constructing a perturbation expansion of the concentration in powers of $\mathrm{Da}$. We then analyze the behaviour of the system at early times and derive explicit expressions for both the concentration and the position of the reacting front in the small-time limit.
@@ 3.1 Perturbation solution for small $\mathrm{Da}$
$+\left.\frac{12\mathrm{T}_{m}^{-2}}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\left(\text{Arctanh}\left(\frac{1-2s}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\right)+\text{Arctanh}\left(\frac{1}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\right)\right)\right]\,.$
@@
$t_{f}=\frac{1}{6}\left[1+\mathrm{Da}\left(1-6\mathrm{T}_{m}^{-1}+\frac{24\mathrm{T}_{m}^{-2}}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\text{Arctanh}\left(\frac{1}{\sqrt{1+4\mathrm{T}_{m}^{-1}}}\right)\right)\right]+\mathrm{T}_{m}^{-1}\,,$ (49)
@@
Similar to the PSS case, an expression for the kinetic-controlled regime can be obtained by assuming $\mathrm{T}_{m}\ll 1$ in the perturbation solution (3.1). However, this gives $s\approx 1-\mathrm{T}_{m}t$ as in the PSS case. In this case, the final time can be approximated as
$t_{f}=\mathrm{T}_{m}^{-1}\,.$ (50)
In the diffusion-controlled limit, $\mathrm{T}_{m}\gg 1$, the interface position can be cast as
$t\approx\frac{1-s^{2}}{2}-\frac{1-s^{3}}{3}+\frac{\mathrm{Da}}{6}(1-s)^{2}\,,$ (51)
while the final time is
$t_{f}=\frac{1}{6}\left(1+\mathrm{Da}\right)\,.$ (52)
@@ 3.2 Small time solution
As discussed in Section 2.2, the PSS model is based on neglecting the time derivative in the diffusion equation, based on the assumption $\mathrm{Da}\ll 1$. If we consider a very early time-scale $\tau=\mathcal{O}(\mathrm{Da})$ or smaller then the PSS model no longer holds. In this section, we develop an approximate analytical solution valid in the small-time regime, enabling us to better understand the system’s initial transient behaviour and provide an initial condition for a numerical solution.
@@
As only a short period of time has elapsed, the interface is not far away from $R^{*}$, the radius of the particle. Therefore, it is useful to propose an ansatz as
@@
This solution coincides with the kinetic-controlled result (36).
@@
An interesting point is that although the PSS does not hold at small times, the small time behaviour matches the kinetic-controlled solution (even in diffusion dominated situations). This means that, even though the kinetic solution does not hold for early times, according to the restrictions of the PSS approximation, it still captures the correct behaviour. This move away from the diffusion form may be inferred from the solution (34) by setting $s=1-\varepsilon f(t)$, where $\varepsilon\ll 1$
$t$ $=\frac{s^{3}-1}{3}-\frac{s^{2}-1}{2}-\mathrm{T}_{m}^{-1}(s-1)=-\frac{3\varepsilon f}{3}+\frac{2\varepsilon f}{2}-\mathrm{T}_{m}^{-1}\varepsilon f+\mathcal{O}({\varepsilon^{2}})\,$ (70)
@@ 4 Numerical solution
The semi-implicit scheme (4) can be formulated as a matrix linear system which can be solved by inverting the matrix of the system at each time-step. All numerical solutions were obtained in a uniform mesh of $10^{3}$ points, using a time-step size of $\Delta t=4\times 10^{-6}$. The implementation was carried out using a custom MATLAB code [40].
@@ 5 Results and discussion
In this section, we first compare the approximate analytical solutions presented in Section 3 with the numerical results from Section 4, focusing on the solution behaviour across different limits of the dimensionless parameters $\mathrm{Da}$ and $\mathrm{T}_{m}$. We then demonstrate how the analytical solutions of the SCM found can be used to fit experimental data for the purpose of estimating unknown physical parameters of the process.
@@ 5.1 Discussion of solutions and limiting behaviours
In typical SCM-based experiments, measurements are used to determine the reacted fraction $X(t)$, which, as shown in (31), depends on the position of the reaction front $s(t)$. Thus, the evolution of $X(t)$ is directly governed by the behaviour of $s(t)$.
In Figure 2 we show the evolution of $s(t)$ for various combinations of $\mathrm{Da}=0.01$, 0.10.10.1, 111, and $\mathrm{T}_{m}=0.1$, 111, 101010, as predicted by the numerical solution, the PSS approximation (34), and the first-order perturbation solution (3.1). The values of $\mathrm{Da}$ are kept below 1, as this range reflects the most physically realistic scenarios, corresponding to typical fluid-to-solid density ratios, as discussed in Section 2.1. Then, the values $\mathrm{T}_{m}=0.1$ to $\mathrm{T}_{m}=10$ are chosen to study the diffusion and kinetic-controlled regimes, respectively, which are the ones typically discussed in the literature.
@@
Focusing on the comparison between the numerical and approximate analytical solutions, we observe that the discrepancies increase with larger values of $\mathrm{Da}$. This behaviour is expected: the PSS solution is derived in the limit $\mathrm{Da}=0$, and the first-order perturbation solution is based on an expansion for $\mathrm{Da}\ll 1$. Therefore, the smaller the value of $\mathrm{Da}$, the closer the approximate solutions are to the full numerical solution, which retains all terms in the model. This trend is especially noticeable in the cases with $\mathrm{T}_{m}=10$ (right panels of Figure 2).
@@
An interesting feature is the distinct change in the behaviour of $s(t)$ around $\mathrm{T}_{m}=1$, which marks the transition between the diffusion-controlled ($\mathrm{T}_{m}\gg 1$) and kinetic-controlled ($\mathrm{T}_{m}\ll 1$) regimes. In particular, the kinetic-controlled regime exhibits a virtually linear evolution of $s(t)$ (left panels of Figure 2), consistent with the linear expressions obtained in (36) and (51) for the PSS and perturbation solutions, respectively, in the limit $\mathrm{T}_{m}\ll 1$. The trends observed for large $\mathrm{T}_{m}$ are characteristic of the classical Stefan problem in spherical geometry (see, e.g., [22, 41, 42]). This is expected, as setting $\mathrm{T}_{m}^{-1}=0$ in equations (26)–(29), previously rewriting (29) in terms of the gradient of $c$ using (28), reduces the SCM to the classical one-phase Stefan problem for a sphere.
@@
The case with $\mathrm{Da}=0.1$, $\mathrm{T}_{m}=0.1$, corresponding to the kinetic-limited regime, is characterized by concentrations remaining close to 1 throughout the domain. This reflects the fact that diffusion is much faster than reaction, allowing the concentration near the moving boundary to rapidly equilibrate with the external concentration due to efficient mass transport. The inset plot for $\mathrm{Da}=0.1$, $\mathrm{T}_{m}=0.1$ provides a close-up view of the concentration profiles near $c\approx 1$, revealing shapes qualitatively similar to those in the diffusion-limited regime, though with concentration variations that are an order of magnitude smaller. Finally, we note the excellent agreement between the numerical solution and the first-order perturbation approximation, with only minor discrepancies appearing near the end of the process in the diffusion-limited case.
@@ 5.2 Comparison to synthetic data
Experimental studies often employ fitting procedures to determine unknown physical parameters of the process. In solid-gas reaction experiments using the SCM, typical unknowns include the reaction rate constant $k$, the diffusion coefficient of the ash layer $D$ or the equilibrium concentration $c_{\rm eq}^{*}$. The experimental data commonly consist of the reacted fraction $X$ measured as a function of time, $t^{*}$. Therefore, the fitting process generally involves estimating the unknown parameters by fitting datasets of the form $(t_{\text{exp},i},X_{\text{exp},i})_{i\in\{1,\dots,N\}}$. In this section, we show how the small time limit (63) and the perturbation solution (3.1) can be used to determine the parameters $k$, $D$ and $c_{\rm eq}^{*}$ from an experimental dataset. In addition, the PSS approximation (34) will also be tested as it is widely used in the literature.
@@
The fitting process is done in two steps. The first, involves finding a reference value of the parameter group $\tau_{\text{eff}}\equiv\tau/\mathrm{T}_{m}$ using the dimensional form of the small time solution (63), i.e. $s(t^{*})=1-(\mathrm{T}_{m}/\tau)\,t^{*}=1-(t^{*}/\tau_{\text{eff}})$. Since the small time limits are short-lived, we take the first two points of the dataset to compute the slope $1/\tau_{\text{eff}}$ of the straight line connecting them. The second, involves using (3.1) to find the values of $\mathrm{T}_{m}$ and $\mathrm{Da}$, that best minimize the quadratic error
@@
In Figure 4, we present the predictions of $s(t^{*})$ and $X(t^{*})$ obtained from the fitting procedure using both the first-order perturbation and PSS solutions, plotted alongside the synthetic data. For the dataset corresponding to $\mathrm{Da}=0.1$, both the first-order perturbation and PSS solutions closely match the data, and the improvement offered by the perturbation solution is not visually evident. In contrast, for the second dataset ($\mathrm{Da}=1$), the first-order perturbation solution provides a visibly better fit than the PSS. This result is expected, since the first-order correction is of order $\mathcal{O}(\mathrm{Da})$; as $\mathrm{Da}$ increases, the correction captures the behaviour of $s(t^{*})$ and $X(t^{*})$ more accurately than the PSS solution, which assumes $\mathrm{Da}=0$.
The last step requires translating the results of the fitting procedure to values of physical process parameters. As mentioned previously, in solid-gas reaction systems, the most common unknowns are $k$, $D$ and $c_{\rm eq}^{*}$. Recalling the definitions of the fitting parameters in terms of the original physical quantities
$\mathrm{Da}=\frac{M_{g}(c_{R}^{*}-c_{\rm eq}^{*})}{\rho_{\beta}},\quad\mathrm{T}_{m}=\frac{R^{*}k}{D},\quad\tau=\mathrm{T}_{m}\,\tau_{\text{eff}}=\frac{\rho_{\beta}R^{*2}}{M_{g}(c_{R}^{*}-c_{\rm eq}^{*})D}\,,$
@@
$c_{\rm eq}^{*}=c_{R}^{*}+\frac{\mathrm{Da}\rho_{\beta}}{M_{g}}\,,\quad D=\frac{R^{*2}}{\mathrm{Da}\mathrm{T}_{m}\,\tau_{\text{eff}}}\,,\quad k=\frac{R^{*}}{\mathrm{Da}\tau_{\text{eff}}}\,,$
if the 1st order perturbation solution is used to fit the data. Although, as demonstrated, the PSS model tends to produce larger errors compared to the first-order correction during fitting, it may still yield approximate values for $k$ and $D$, provided that $c_{\rm eq}^{*}$ is either known or negligible relative to $c_{R}^{*}$. In such cases, $c_{\rm eq}^{*}$ can be omitted from the expressions for $\tau$ and $\mathrm{Da}$.